Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
P.O. Box 117000
Gainesville, FL 32611-7000

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
November 8, 2011

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Michael Jay, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan Wack
Announcements
  1. Open Enrollment November 7-18.
  2. Judith Roberts orientation - November 17, she will be at ECOB to discuss IT training needs and to become familiar with our department.
Ongoing IT Updates

Hardware
  1. App streaming - testing DLC and Cataloging software.
  2. Migration - UAC is being turned off of public machines. West Circulation, Cataloging and Administration are next on the list.
  3. 1A - wall behind the podium has been completely gutted and redone. All cables, networking and equipment will be locked up in the cabinet. The phone will be accessible. AVI should be returning on Thursday to finish up the job. Some new equipment may be purchased for 1A by Academic Technology so they can set up all of the necessary equipment when Elias speaks there.
  4. VDI Project - thin clients are set up in ECOB and West shops. Benchmark testing software installed by Logan. VMware View reinstalled.
  5. Images - still planning to remove language packs from all public images. Reboot testing - networking issue in MSL. Will brooch the topic again with Branch Chairs. Trish has been testing the new staff image with log off, .Net framework and updates. Need to recreate Netbook images when Chris' software is done.
  6. Smart Control - NIC issues - Josh has been on the phone with the company.
  7. Privilege Authority policy - up and running. We will push it so that staff can update the following software: Java, Iliad, AdobeAIR/Flash/Shockwave, GoToMeeting, Eliminate, WebEx, ITunes, QuickTime, VLC, OCLC, UPS World Ship, Smart Board Software
  8. West/MSL wireless network changes are in place. Need to reimage netbooks.
  9. Printing changes - 64 bit drivers for public printing in the works. Have contacted Canon about Mac drivers. Code for staff departmental printers for printing confidential information.
  10. Symantec PGP encryption - will download and get in place.
  11. Way-Finding system - digital signage was not what Judy wanted. She is interested in way-finding signage: Nano Nation interactive boards.
Software
  1. Document disposition database - Tom finished the last request. It has been rolled out. Done.
  2. Netbook NIC enabler - Chris is working on it.
  3. Kuale OLE - Chris, Rachel and Judy are attending Kuale Days.
  4. DLC - meeting with Lois Widmer. DLC looking into new version of Web.
  5. DLC-to-Hathi Trust - move to pending.
  6. Film Tracker - back burner, over taken by other projects.
  7. Storage Gap Tracker - Chris is working on it.
  8. Macros - three new ones.
  9. ASERL Gap Analysis - called Lyricist, they are sending all the stuff for tools.
  10. ARES-Sakai authentication - Michael is working with Warren Curry on this - it's close to done.
  11. Graham papers data conversion - new request. Need to convert an old database into a readable format.
  12. Circulation switch to "Library Guest" accounts - they will no longer be issuing general Library accounts. Instead, they will create Library Guest accounts that will allow for repeated use and be slightly more permanent.
Network/Server Administration
  1. New OLE server up and running. Chris Case was having an issue with it, but that has been resolved.
  2. No word on new departmental VPN.
  3. ALADN - no movement.
  4. GitHub server is working - basic install. Runs Git and GitWeb. Mark and Winston will meet and go over.
  5. Finishing up moving the blog server.
  6. PeopleCounter server messed up - meeting to discuss. Once this final issue is resolved, it will be done.
  7. Ports dropping out in the ECOB shop - ticket generated with CNS.